Two cars leave an intersection at the same time. One drives east while the other drives south at 10 miles per hour faster than the other. After 2 hours, the cars are 100 miles apart. How fast is the southbound car traveling?
Answer by rwm(914)

We are dealing with a right triangle here and the 100 miles is the hypotenuse.
(e*2)^2+((e+10)*(2))^2=100^2
8x^2+80x+400 = 10000
8 x^2+80 x-9600 = 0
e=-40
e=30 mph
south is 30+10=40 mph
we disregard negative answers since we can't have a negative speed.